The length of a rectangular flower garden is 5 feet longer than its width. If the area of the garden is 50 square feet, find the length of the garden.

thank you!

Answer by checkley79(3341)

AREA=LW
L=W+5
50=(W+5)W
50=W^2+5W
W^2+5W-50=0
(W+10)(W-5)=0
W-5=0
W=5 ANS. FOR THE WIDTH.
L=5+5=10 ANS. FOR THE LENGTH.
PROOF:
50=10*5
50=50